the equation of the line that has a gradient of 3 and passes through the y-axis at (0,1)


1
Expert's answer
2021-07-16T01:23:26-0400

"Gradient =\frac {\Delta Y} {\Delta X}"


"3=\frac {y-1}{x-0}"


"3x=y-1"


Putting the equation in the form y=mx+c


"y=3x+1"